Define permutation function and find out all permutation on the set S = {π‘Ž, 𝑏, 𝑐}

1
Expert's answer
2022-03-22T18:29:38-0400

"P\left( S \right) =\left| S \right|!=3!=6\\All\,\,permutations:\\\left( a,b,c \right) ,\left( a,c,b \right) ,\left( b,a,c \right) ,\left( b,c,a \right) ,\left( c,a,b \right) ,\left( c,b,a \right)"